Grade 11Magical Mathematics[Interesting Approach]

John ordered 4 pairs of black socks and some additional pairs of blue socks. The price of the black socks per pair was twice that of the blue. When the order was filled,it was found that the no. Of pairs of the two colours had been interchanged. This increased the bill by 50%. What was the ratio of the no. Of pairs of black socks to the no. Of pairs of blue socks in the original order? Please provide the solution.

Shrey
12 Years agoGrade 11

1 Answer

sunil raikwar
ApprovedApproved Tutor Answer12 Years ago
Let the price of a blue socks pair be Rs. a then price of a black pair=Rs 2a & John ordered n pair of blue socks according to question correct bill=(4*2a+n*a) Incorrect bill=n*2a+4*a Second bill is 50% more than the first bill 150%of(4*2a+n*a)=n*2a+4*a 24a+3na=4na+8a 16a=na n=16 Thanks & Regards Sunil Raikwar askIITians faculty